OpenStudy (needshelpdfdf):

Solve for x. 2x2 − 4x = 0

OpenStudy (eliesaab):

\[ 2 x^2 -4 x=2x(x-2)=0\]

OpenStudy (eliesaab):

The answer is?

OpenStudy (eliesaab):

\[ 2 x=0 \implies x=0\\ x-2=0\implies x=2 \]

OpenStudy (needshelpdfdf):

0,-2?

OpenStudy (eliesaab):

No, look at my second relpy

OpenStudy (needshelpdfdf):

oh 2,4?

OpenStudy (eliesaab):

third reply

OpenStudy (needshelpdfdf):

oh I see third reply

OpenStudy (eliesaab):

0 is fine, the second one is x=2

OpenStudy (needshelpdfdf):

oh 0,2

OpenStudy (eliesaab):

Yes

OpenStudy (needshelpdfdf):

I feel stupid

OpenStudy (eliesaab):

With repetition, you get better